# HOPSE: Scalable Higher-Order Positional and Structural Encoder for Combinatorial Representations

### v1

@@ F.5.1 Analysis of performance across neighborhoods
Graph Tasks Topological Tasks Nbhd MUTAG ($\uparrow$) PROTEINS ($\uparrow$) NCI1 ($\uparrow$) NCI109 ($\uparrow$) ZINC ($\downarrow$) NAME ($\uparrow$) ORIENT ($\uparrow$) $\beta_{1}$ ($\uparrow$) $\beta_{2}$($\uparrow$) HOPSE-M Adj-1 86.38 $\pm$ 6.31 74.70 $\pm$ 4.04 77.35 $\pm$ 0.81 75.41 $\pm$ 1.17 0.40 $\pm$ 0.01 79.58 $\pm$ 2.88 60.93 $\pm$ 3.03 88.35 $\pm$ 0.09 28.25 $\pm$ 2.71 Adj-2 85.53 $\pm$ 7.12 73.19 $\pm$ 3.32 76.26 $\pm$ 1.59 75.59 $\pm$ 0.41 0.39 $\pm$ 0.02 80.35 $\pm$ 3.96 74.63 $\pm$ 1.77 89.97 $\pm$ 0.18 64.06 $\pm$ 2.27 Adj-3 85.96 $\pm$ 5.75 75.13 $\pm$ 3.66 76.44 $\pm$ 1.44 75.22 $\pm$ 2.06 0.39 $\pm$ 0.02 84.72 $\pm$ 1.23 78.19 $\pm$ 2.70 90.13 $\pm$ 0.49 70.10 $\pm$ 1.86 Inc-1 85.11 $\pm$ 6.20 73.98 $\pm$ 2.97 76.40 $\pm$ 0.99 76.09 $\pm$ 1.50 0.40 $\pm$ 0.01 84.01 $\pm$ 6.84 65.10 $\pm$ 9.66 90.06 $\pm$ 0.09 65.02 $\pm$ 1.58 Inc-2 84.26 $\pm$ 6.13 73.69 $\pm$ 2.76 76.40 $\pm$ 1.45 75.39 $\pm$ 1.13 0.42 $\pm$ 0.05 80.67 $\pm$ 2.88 69.51 $\pm$ 6.86 89.51 $\pm$ 0.94 64.28 $\pm$ 3.08 Inc-3 85.11 $\pm$ 5.83 73.55 $\pm$ 3.38 76.56 $\pm$ 1.40 76.34 $\pm$ 1.40 0.42 $\pm$ 0.02 82.56 $\pm$ 4.15 77.35 $\pm$ 2.12 89.89 $\pm$ 0.13 63.12 $\pm$ 1.68 Mix-1 84.26 $\pm$ 5.12 73.69 $\pm$ 3.21 76.44 $\pm$ 1.62 75.57 $\pm$ 1.09 0.40 $\pm$ 0.01 90.37 $\pm$ 1.29 72.64 $\pm$ 7.53 90.21 $\pm$ 0.18 67.14 $\pm$ 2.18 Mix-2 83.83 $\pm$ 6.49 73.41 $\pm$ 2.95 75.14 $\pm$ 1.59 74.87 $\pm$ 0.94 0.41 $\pm$ 0.02 91.50 $\pm$ 1.45 80.68 $\pm$ 1.72 90.26 $\pm$ 0.55 71.69 $\pm$ 1.50 HOPSE-G Adj-1 82.98 $\pm$ 7.22 74.05 $\pm$ 2.17 77.18 $\pm$ 1.45 76.21 $\pm$ 0.65 0.44 $\pm$ 0.01 80.43 $\pm$ 1.03 60.92 $\pm$ 1.18 88.18 $\pm$ 0.03 24.59 $\pm$ 1.87 Adj-2 85.53 $\pm$ 2.77 74.05 $\pm$ 3.08 77.51 $\pm$ 1.15 74.91 $\pm$ 1.17 0.43 $\pm$ 0.00 80.19 $\pm$ 1.13 61.88 $\pm$ 0.83 88.13 $\pm$ 0.02 20.31 $\pm$ 2.57 Adj-3 85.11 $\pm$ 3.36 73.26 $\pm$ 4.20 77.32 $\pm$ 1.19 74.79 $\pm$ 0.91 0.43 $\pm$ 0.01 79.11 $\pm$ 1.53 61.86 $\pm$ 1.10 88.11 $\pm$ 0.02 19.08 $\pm$ 3.57 Inc-1 84.68 $\pm$ 3.50 73.19 $\pm$ 3.52 77.70 $\pm$ 1.40 76.36 $\pm$ 1.75 0.44 $\pm$ 0.00 80.70 $\pm$ 0.80 62.17 $\pm$ 0.98 87.31 $\pm$ 1.17 20.47 $\pm$ 3.22 Inc-2 82.13 $\pm$ 5.75 76.27 $\pm$ 2.10 75.66 $\pm$ 1.52 74.17 $\pm$ 1.99 0.45 $\pm$ 0.02 78.92 $\pm$ 1.25 59.57 $\pm$ 1.56 88.08 $\pm$ 0.08 20.28 $\pm$ 3.48 Inc-3 77.87 $\pm$ 4.90 75.48 $\pm$ 3.15 76.15 $\pm$ 1.77 75.14 $\pm$ 0.75 0.44 $\pm$ 0.01 78.78 $\pm$ 0.74 60.35 $\pm$ 1.05 88.12 $\pm$ 0.04 18.82 $\pm$ 3.54 Mix-1 79.15 $\pm$ 7.88 75.70 $\pm$ 3.24 76.50 $\pm$ 1.09 75.35 $\pm$ 0.97 0.45 $\pm$ 0.01 79.76 $\pm$ 0.51 60.33 $\pm$ 1.60 88.07 $\pm$ 0.09 23.48 $\pm$ 3.44 Mix-2 83.40 $\pm$ 8.56 74.05 $\pm$ 3.39 77.02 $\pm$ 0.92 74.75 $\pm$ 1.47 0.46 $\pm$ 0.01 81.75 $\pm$ 1.26 61.32 $\pm$ 2.67 88.28 $\pm$ 0.08 35.37 $\pm$ 2.25
Efficacy of neighborhoods by domain.
One important observation is how neighborhood efficacy varies between the simplicial and cell domains. While in the simplicial domain, adding mixed neighborhoods occasionally leads to performance degradation for some datasets (e.g., PROTEINS where Inc-1 outperforms Mix configurations for HOPSE-M), in the cell domain, more comprehensive neighborhood sets either improve or do not significantly impact performance. This domain-dependent behavior suggests that the optimal neighborhood configuration should be ideally selected based on task complexity and the topological representation chosen for the data; however, choosing comprehensive neighborhood sets on average allows for achieving good performance while not harming the computational cost (see Section F.5.2 for details on computational analysis). These findings demonstrate that the choice of neighborhood configuration represents a critical design decision in topological learning frameworks. While simpler configurations may suffice for some graph benchmarks, an effective strategy is to consider a greater neighbourhood set ${\mathcal{N}}_{C}$ to capture the full expressive power of topological structures. It is especially evident on complex tasks, topological benchmarking datasets (MANTRA), and relatively large datasets like ZINC.
Table 7: Performance per neighborhood on cell. Results are shown as mean $\pm$ standard deviation. The best result per domain and model is in bold and results within one standard deviation of the lowest per domain are highlighted in blue. In smaller and simpler datasets such as MUTAG, PROTEINS, NCI1, and NCI109, the best neighborhoods are not always the most encompassing, but they perform closely to others. However, on a more complex dataset such as ZINC there is a significant difference between simpler neighborhoods like the Adj[1-3] and Mix-2, particularly in the case of HOPSE-G.

### v2

@@ Appendix H Performance ablation across different neighborhoods
Graph Tasks Topological Tasks $\mathcal{N}_{S}$ MUTAG ($\uparrow$) PROTEINS ($\uparrow$) NCI1 ($\uparrow$) NCI109 ($\uparrow$) ZINC ($\downarrow$) NAME ($\uparrow$) ORIENT ($\uparrow$) $\beta_{1}$ ($\uparrow$) $\beta_{2}$($\uparrow$) HOPSE-M Adj-1 $83.40\pm 0.07$.12) $76.13\pm 0.03$.09) $75.97\pm 0.01$.18) $75.70\pm 0.01$.17) $0.43\pm 0.00$.01) $85.36\pm 0.08$.46) $83.76\pm 0.00$.94) $90.32\pm 0.00$.09) $68.79\pm 0.00$.78) Adj-2 $85.96\pm 0.05$.55) $76.85\pm 0.01$.75) $76.48\pm 0.01$.44) $75.37\pm 0.00$.98) $0.42\pm 0.00$.03) $94.45\pm 0.00$.82) $89.24\pm 0.00$.22) $91.72\pm 0.01$.17) $78.45\pm 0.01$.92) Adj-3 $84.68\pm 0.05$.51) $76.13\pm 0.03$.01) $75.64\pm 0.00$.76) $75.06\pm 0.00$.84) $0.43\pm 0.00$.01) $96.01\pm 0.00$.54) $88.60\pm 0.00$.77) $91.13\pm 0.01$.20) $79.12\pm 0.01$.11) Inc-1 $84.26\pm 0.05$.55) $76.06\pm 0.03$.05) $75.51\pm 0.01$.03) $75.20\pm 0.00$.93) $0.42\pm 0.00$.01) $96.27\pm 0.00$.28) $87.34\pm 0.00$.74) $90.76\pm 0.01$.57) $78.32\pm 0.01$.13) Inc-2 $85.11\pm 0.07$.22) $75.27\pm 0.01$.88) $75.51\pm 0.00$.53) $74.42\pm 0.00$.80) $0.43\pm 0.00$.02) $95.96\pm 0.00$.29) $86.35\pm 0.02$.15) $91.22\pm 0.00$.06) $79.08\pm 0.00$.75) Inc-3 $84.68\pm 0.06$.96) $76.13\pm 0.01$.84) $76.01\pm 0.00$.64) $74.77\pm 0.01$.21) $0.42\pm 0.00$.01) $96.07\pm 0.00$.19) $85.74\pm 0.01$.03) $91.07\pm 0.00$.10) $77.73\pm 0.01$.69) Mix-1 $82.98\pm 0.03$.98) $76.56\pm 0.02$.84) $75.66\pm 0.00$.32) $75.84\pm 0.00$.50) $0.43\pm 0.00$.01) $93.84\pm 0.01$.76) $87.42\pm 0.00$.64) $91.63\pm 0.00$.93) $80.45\pm 0.01$.06) Mix-2 $83.83\pm 0.04$.90) $76.06\pm 0.02$.31) $75.21\pm 0.01$.30) $75.90\pm 0.00$.48) $0.44\pm 0.00$.01) $95.71\pm 0.01$.41) $89.23\pm 0.00$.15) $90.87\pm 0.01$.07) $82.29\pm 0.00$.46) HOPSE-G Adj-1 $82.98\pm 0.07$.22) $74.05\pm 0.02$.17) $77.18\pm 0.01$.45) $76.21\pm 0.00$.65) $0.44\pm 0.00$.01) $80.43\pm 0.01$.03) $60.92\pm 0.01$.18) $88.18\pm 0.00$.03) $24.59\pm 0.01$.87) Adj-2 $85.53\pm 0.02$.77) $74.05\pm 0.03$.08) $77.51\pm 0.01$.15) $74.91\pm 0.01$.17) $43\pm 0.00.00$) $80.19\pm 0.01$.13) $61.88\pm 0.00$.83) $88.13\pm 0.00$.02) $20.31\pm 0.02$.57) Adj-3 $85.11\pm 0.03$.36) $73.26\pm 0.04$.20) $77.32\pm 0.01$.19) $74.79\pm 0.00$.91) $0.43\pm 0.00$.01) $79.11\pm 0.01$.53) $61.86\pm 0.01$.10) $88.11\pm 0.00$.02) $19.08\pm 0.03$.57) Inc-1 $84.68\pm 0.03$.50) $73.19\pm 0.03$.52) $77.70\pm 0.01$.40) $76.36\pm 0.01$.75) $44\pm 0.00.00$) $80.70\pm 0.00$.80) $62.17\pm 0.00$.98) $87.31\pm 0.01$.17) $0.47\pm 0.03.22$) Inc-2 $82.13\pm 0.05$.75) $76.27\pm 0.02$.10) $75.66\pm 0.01$.52) $74.17\pm 0.01$.99) $0.45\pm 0.00$.02) $78.92\pm 0.01$.25) $59.57\pm 0.01$.56) $88.08\pm 0.00$.08) $20.28\pm 0.03$.48) Inc-3 $77.87\pm 0.04$.90) $75.48\pm 0.03$.15) $6.15\pm 0.01.77$) $75.14\pm 0.00$.75) $0.44\pm 0.00$.01) $78.78\pm 0.00$.74) $60.35\pm 0.01$.05) $88.12\pm 0.00$.04) $18.82\pm 0.03$.54) Mix-1 $79.15\pm 0.07$.88) $75.70\pm 0.03$.24) $76.50\pm 0.01$.09) $75.35\pm 0.00$.97) $0.45\pm 0.00$.01) $79.76\pm 0.00$.51) $60.33\pm 0.01$.60) $88.07\pm 0.00$.09) $23.48\pm 0.03$.44) Mix-2 $83.40\pm 0.08$.56) $74.05\pm 0.03$.39) $77.02\pm 0.00$.92) $74.75\pm 0.01$.47) $0.46\pm 0.00$.01) $81.75\pm 0.01$.26) $61.32\pm 0.02$.67) $88.28\pm 0.00$.08) $35.37\pm 0.02$.25)
Table 7: Performance ablation of $\mathcal{N}_{S}$ on cell domain.